What is the effective annual rate of 10% compounded monthly?

1. 9.27%

2. 10.00%

3. 10.25%

4. 10.38%

5. 10.47%
Answer #1

Annual Rate = 10%

Interest rate per period = 10% / 12

= 0.8333%

Effective Annual rate = [( 1 + Interest rate per period) ^ Number of Compounding] - 1

= ( 1 + 0.8333%)^12 - 1

= ( 1.0083333) ^ 12 - 1

= 1.1047126 - 1

= 0.104712 or 10.4712%

Rounded to two decimals = 10.47% [ Option E is the correct answer]